OpenStudy (anonymous):

I fan and give medals!!! A solution of salt and water contains 70 grams of water per 140 milliliters of the solution. If 1 mole of water weighs 16 grams, _____ moles of water would be present in 35 milliliters of the solution. [Round off your answer to two decimal places]

OpenStudy (anonymous):

Fractions. 70 over 140 = x over 35. 140/35=4 Then, 4*70=280 x=280. 280/16=17.5 moles The answer should be 17.5 moles.